OpenStudy (anonymous):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.4 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 9 cm.

OpenStudy (anonymous):

V=\[\frac{4}{3} \pi \left(\frac{D}{2}\right)^3\] \[V=\frac{D^3 \pi }{6}\] \[\frac{\text{dV}}{\text{dt}}=\frac{3\pi }{6}D^2 \frac{\text{dD}}{\text{dt}}\] \[\frac{\text{dv}}{\text{dt}}=\frac{\pi }{2}D^2 \frac{\text{dD}}{\text{dt}}\] we know dD/dt is -0.4 D=9 So let's plug them in \[\frac{\text{dv}}{\text{dt}}=\frac{\pi }{2}9^2 -.04\]=-5.08938

OpenStudy (anonymous):

I appreciate the time you've taken to elaborately explain this, but are you sure about your answer?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

It's telling me that it's incorrect, and I don't think my answer was anywhere near that either (though chances are that I'm doing it wrong since I'm here asking for help)

OpenStudy (anonymous):

Do you have the answer?

OpenStudy (anonymous):

I'm working on 8 similar problems simultaneously, so I can't remember which answer pertains to which problem :/

OpenStudy (anonymous):

so, we have no way to verfiy what the answer is?

OpenStudy (anonymous):

maybe it want exact answer?

OpenStudy (anonymous):

we do, I plug it in to my retarded online thing and it tells me whether it's right or wrong.

OpenStudy (anonymous):

Cause what i am thinking is, that the differntiation of the formula for the volume of a sphere might be wrong. I think it is something like this: \[dv/dt=3(4/3)\pi(d/2)^2\]

OpenStudy (anonymous):

I fortgot to multplu the right hand side by *dd/dt

OpenStudy (anonymous):

so something like this: dv/dt=4pi(d/2)^2*dd/dt, where dd/dt is the change in diamter

OpenStudy (anonymous):

Then we substitue the know values into that equation

OpenStudy (anonymous):

so we get, dv/dt=(4)(pi)(4.5)^2(0.4)

OpenStudy (anonymous):

I found pretty much the same problem online, followed the same method and still couldn't get the answer. The first one is the exact same problem, yet apparently wrong answer. http://answers.yahoo.com/question/index?qid=20101108131052AAi84k0 http://answers.yahoo.com/question/index?qid=20090225144157AA7JlEO

OpenStudy (anonymous):

what program are you tyrp typing it in to check for the answer?

OpenStudy (anonymous):

it's an online homework program which prompts me for my answer; it's called webwork. Most of the errors are due to decimals, but not with these.

OpenStudy (anonymous):

Is the answer 50.893 or 101.787

OpenStudy (anonymous):

the first, omg! how did you get it?

OpenStudy (anonymous):

I am a beast man, what can i say

OpenStudy (anonymous):

lol just kiddin

OpenStudy (anonymous):

Okay well here it is, the first guy that answered differentiatied the volume of the sphere wrong.

OpenStudy (anonymous):

The volume of the sphere should diffentiate to : dv/dt=4*pi*(d/2)^2*dd/dt

OpenStudy (anonymous):

I was wondering why that was chosen as best answer...

OpenStudy (anonymous):

Now, you see where i have d/2, that is basically saying or means radius. But now since i have chosen to work with the radius, i have cut vaulue given in the problem in half, that is to say that i want to work with 0.2 not 0.4( for dd.dt or the rate at which the diameter is decreasing). So i simply susbtitute my numbers into the deferentiatied eqaution.

OpenStudy (anonymous):

Then after multiplying things out we get: dv/dt=(4)(pi)(20.25)(0.2)=50.893

OpenStudy (anonymous):

lol, I see. Thanks man

OpenStudy (anonymous):

no prob.

OpenStudy (anonymous):

My differentiation was right but I plug in wrong number dD/dt=.4 D=9\[\frac{\pi }{2}9^2 (.4)\] =50.8938